### v2

# FLoRA: An Advanced AI-Powered Engine to Facilitate Hybrid Human-AI Regulated Learning

Abstract
SRL, defined as learners’ ability to systematically plan, monitor, and regulate their learning activities, is crucial for sustained academic achievement and lifelong learning competencies. Emerging Artificial Intelligence (AI) developments profoundly influence SRL interactions by potentially either diminishing or strengthening learners’ opportunities to exercise their own regulatory skills. Recent literature emphasizes a balanced approach termed Hybrid Human-AI Regulated Learning (HHAIRL), in which AI provides targeted, timely scaffolding while preserving the learners’ role as active decision-makers and reflective monitors of their learning process. Nevertheless, existing digital tools frequently fall short, lacking adaptability, focusing narrowly on isolated SRL phases, and insufficiently support meaningful human-AI interactions. In response, this paper introduces the enhanced FLoRA Engine, which incorporates advanced Generative Artificial Intelligence (GenAI) features and state-of-the-art learning analytics, explicitly grounded in SRL and HHAIRL theories. The FLoRA Engine offers instrumentation tools such as collaborative writing, multi-agents chatbot, and detailed learning trace logging to support dynamic, adaptive scaffolding tailored to individual needs in real time. We further present a summary of several research studies that provide the validations for and illustrate how these instrumentation tools can be utilized in real-world educational and experimental contexts. These studies demonstrate the effectiveness of FLoRA Engine in fostering SRL and HHAIRL, providing both theoretical insights and practical solutions for the future of AI-enhanced learning context.
keywords:
[…]
{highlights}
The FLoRA Engine provides advanced instrumentation tools to effectively detect, measure, and facilitate self-regulated learning (SRL) processes.
Leveraging GenAI and learning analytics, FLoRA Engine delivers personalized scaffolds that dynamically support diverse learners throughout the SRL cycle.
The FLoRA Engine exemplifies the potential of AI-enhanced platforms to address the complexities of hybrid human–AI regulation in SRL, as demonstrated across multiple educational case studies.
1 Introduction
Self-regulated learning (SRL) is often defined as a learner’s ability to plan, monitor, and adjust personal learning processes in order to meet self-determined goals (Azevedo and Wiedbusch, 2023). It encompasses an array of cognitive and metacognitive strategies – such as goal setting, strategic decision-making, and self-monitoring – that together help learners direct their own progress (Alvarez et al., 2022). Through the active exercise of these strategies, learners can deepen their understanding, build resilience, and enhance their capacity to transfer acquired knowledge and skills across contexts (Winne, 2017b). Moreover, SRL is not only pivotal for achieving higher academic performance – extensive research indicates that well-regulated learners generally achieve better outcomes (Dent and Koenka, 2016; Xu et al., 2023; Broadbent and Poon, 2015) – but also for lifelong learning, which has gained heightened importance in an era where constant skill adaptation is a necessity (Seufert, 2018).
Generative Artificial Intelligence (GenAI) and other recent AI advances are reshaping how educators, researchers, and learners view SRL (Järvelä et al., 2023). On one hand, many adaptive learning technologies can offload tasks like monitoring and selecting new materials, which can inadvertently reduce learners’ opportunities to exercise self-regulation (Molenaar et al., 2023b). On the other hand, the integration of data-driven AI support with explicit scaffolds aimed at promoting and reinforcing learners’ own regulatory skills (Molenaar, 2022a; Lim et al., 2024). The AI can step in with timely, data-driven insights while still encouraging learners to make key decisions and reflect on their learning process (Molenaar and Wise, 2022). This synergy leverages the strengths of both AI and the human learner, ensuring that the learner remains at the centre and continues to build the metacognitive monitoring and control skills essential for future learning, with the process termed as Hybrid Human-AI Regulated Learning (HHAIRL) (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022). Still, there is a risk that an excessive reliance on AI could diminish learners’ engagement in important self-regulatory activities (Zhang et al., 2024), such as monitoring their own progress or planning their learning, thereby impeding the growth of robust SRL skills (Authors, 2024a).
Despite increasing attention to Hybrid Human-AI Regulation (Huang, 2024; Nguyen, 2025; Järvelä et al., 2023; Molenaar and Wise, 2022), several gaps and challenges remain. First, existing technologies for the real-time identification of a learner’s SRL are relatively underdeveloped. Many current tools lack adaptability; they often work well for specific tasks but do not account for the diverse realities of educational environments, where students engage with varied materials and may frequently learn collaboratively (Alvarez et al., 2022). In addition, many tools focus on particular SRL phases rather than encompassing the entire cycle of setting goals, monitoring progress, and evaluating outcomes (Alvarez et al., 2022). The design of tools like dashboards that help learners understand and exert control alongside the AI is still at an early stage; simply showing performance metrics is not enough (Matcha et al., 2020). Furthermore, the emergence of AI especially GenAI, capable of providing human-like interactions and demonstrations of expert-level reasoning, significantly alters typical human-tool interaction patterns (Edwards et al., 2025; Kasneci et al., 2023). Learners increasingly trust and rely on GenAI systems due to their sophisticated capacities and human-like interactions, fundamentally changing how learners engage in SRL processes such as goal-setting, monitoring, and reflection (Hauske and Bendel, 2024; Lai, 2024). Nonetheless, existing SRL support tools, primarily designed prior to these advancements, are insufficiently equipped with GenAI-based functionalities to reliably track, interpret, and scaffold these complex human-GenAI interactions (Edwards et al., 2025; Holmes et al., 2023). Finally, many SRL support tools often exhibit strong dependencies on their specific learning contexts (Wong et al., 2019; Matcha et al., 2020), limiting their applicability across diverse disciplines and making them unsuitable for evaluating the effectiveness of HHAIRL in diversified learning scenarios (Molenaar, 2022a).
In response to these limitations, the FLoRA Engine has been developed based on the previous work (Authors, 2025a). Building upon prior version, the current version of FLoRA not only significantly expands the three main modules – instrumentation tools, trace parser, and scaffolding, but also introduces several innovative features to support HHAIRL research. One particularly significant advancement in this iteration is the integration of GenAI technologies, such as GPT-4o, which enable more dynamic, personalised, and adaptive learner-centred support. In parallel, FLoRA incorporates advanced learning analytics methodologies (Gašević et al., 2015; Matcha et al., 2019) grounded firmly in established SRL theories (Saint et al., 2020b; Siadaty et al., 2016), ensuring that insights generated from learner-interaction data remain theoretically coherent and practically meaningful. By aligning closely with emerging trends in AI-enhanced educational practice, FLoRA serves a dual purpose: on one hand, it provides researchers with a comprehensive, robust platform for systematically investigating the complex dynamics of Hybrid Human-AI regulation; on the other, it offers educators and learning scientists a practical example illustrating how GenAI technologies can effectively enhance SRL-supportive educational contexts.
In the following sections, a detailed description of FLoRA Engine’s features and design rationale are presented. Subsequently, empirical evidence gleaned from illustrative case studies are described, clearly demonstrating how these novel features concretely support learners’ SRL capabilities, across diverse, AI-mediated educational scenarios.
2 Theoretical Foundations
2.1 Self-Regulated Learning
Self-regulated learning (SRL) is a comprehensive framework that encompasses the cognitive, metacognitive, behavioural, motivational, and emotional aspects of learning (Zimmerman, 1986; Winne, 2018; Winne and Hadwin, 1998; Winne, 2017a). Through a range of strategies, such as planning, goal setting, self-monitoring, self-instruction, and self-evaluation, SRL enable learners to take control of their own learning processes. By engaging in SRL, learners become proactive in their education, employing various tactics to understand and master new information, while also managing their emotions and staying motivated throughout the learning journey (Panadero, 2017).
The importance of SRL lies in its significant impact on educational outcomes and lifelong learning (Dent and Koenka, 2016; Xu et al., 2023; Broadbent and Poon, 2015; Guo, 2022). SRL equips learners with the skills necessary to adapt to various learning contexts and challenges, fostering autonomy and resilience. When students practice SRL, they are more likely to exhibit higher levels of self-efficacy, persistence, and academic achievement (Uzir et al., 2020; van der Graaf et al., 2022). Moreover, SRL promotes deeper engagement with learning materials (Li and Lajoie, 2022), critical thinking (Anwar and Muti’ah, 2022), and the ability to transfer knowledge to new situations (Stebner et al., 2022; Schuster et al., 2020). As educational environments become increasingly complex, the ability to self-regulate is essential for success. Educators recognise that teaching students how to learn, rather than just what to learn, empowers them to become effective, lifelong learners who can navigate and thrive in a rapidly changing world (Guo, 2022).
2.2 The Detection and Measurement of SRL
Detecting and measuring SRL is pivotal for understanding how learners control and direct their learning processes, especially in computer-based environments (Molenaar et al., 2023a). SRL involves a spectrum of cognitive, metacognitive, motivational and affective processes (Azevedo, 2015; Winne, 2023), such as setting goals, monitoring progress, and adjusting strategies to achieve learning objectives. However, capturing these internal processes poses significant challenges because they are often not directly observable and may not always manifest in overt behaviours that can be easily recorded or quantified (Saint et al., 2020a).
One effective approach to detecting SRL is the utilization of log data and learning analytics within digital learning platforms (Siadaty et al., 2016; Winne, 2022; Bernacki, 2025). As learners interact with computer-based educational resources, they generate extensive trace data—clickstreams, navigation paths, and tool usage logs—that can be analysed to infer SRL activities (Papamitsiou and Economides, 2014; Bannert et al., 2014; Azevedo and Gašević, 2019; Li et al., 2020; Deininger et al., 2025). For instance, if a learner chooses to attempt a quiz before engaging with reading materials, this might indicate proactive planning and goal-setting behaviours. However, not all SRL actions leave discernible traces in the log data, and some SRL processes, particularly metacognitive ones, may remain hidden or indistinct, making it challenging to achieve a comprehensive understanding solely through raw interaction data (Winne, 2017a, 2018).
To address limitations in trace-based measurements of SRL, researchers have developed and integrated instrumentation tools within learning environments that are specifically designed to elicit and capture SRL processes (Winne et al., 2019; van der Graaf et al., 2021). These tools serve a dual purpose: they facilitate learning by providing functionalities that support SRL, and they generate detailed trace data that reflect learners’ cognitive and metacognitive activities. For example, tools that allow learners to highlight text, take notes, or set reminders encourage them to engage in organization, elaboration, and planning—key components of SRL (Paans et al., 2019; Cerezo et al., 2020). A pioneering exemplar of such a learning technology is nStudy (Winne et al., 2019), which offers a suite of features enabling learners to interact deeply with content while simultaneously logging these interactions for analysis. From the learners’ perspective, these tools empower them to actively engage in self-regulation by making their cognitive strategies explicit. For educators and researchers, the instrumentation tools provide valuable data that can be analysed to gain insights into the learners’ SRL processes.
Analysing the rich trace data generated requires sophisticated methods to accurately map observable behaviours to underlying SRL processes (Bannert, 2007; Molenaar et al., 2023a). Conceptualizing SRL as a series of events with increasing complexity—occurrence, contingency, and patterned contingency—provides a structured framework for this analysis (Lim et al., 2021). The occurrence level captures the frequency of specific learning actions, such as the number of times a learner highlights text. The contingency level examines sequences of actions, providing context by looking at how one action follows another—for example, highlighting text immediately after taking a quiz. The patterned contingency level involves identifying recurring sequences or patterns of actions that suggest a strategic approach to learning, such as a consistent cycle of reading, note-taking, and self-testing. By employing temporal analytical methods like process mining and Markov models, researchers can uncover these patterns within the data, revealing how learners self-regulate over time (Saint et al., 2022).
2.3 Hybrid Human-AI and Socially Shared Regulated Learning
Socially-Shared Regulation of Learning (SSRL) is the process through which members of a group collectively plan, monitor and adapt their activity in order to reach a shared learning goal (Järvelä et al., 2018; Haataja et al., 2022). Regulation decisions—clarifying the task, setting sub-goals, checking mutual understanding or reallocating effort—are negotiated and enacted together, not left to any single individual (Zheng et al., 2019; Kent and Cukurova, 2020). Traditionally, teachers have scaffolded these collective metacognitive moves; now GenAI can join the conversation (Järvelä et al., 2023). By analysing multimodal data streams in real time, an AI agent can detect moments when the group drifts apart, highlight emerging misconceptions or surface silent contributions, effectively acting as an additional “group member” that nudges the team toward productive dialogue (Edwards et al., 2025). Thus, AI becomes an enabler of richer SSRL, augmenting human awareness rather than replacing it.
Hybrid Human-AI Regulation (HHAIR) generalises this idea by treating regulation itself as a shared responsibility that can fluidly shift between learners and intelligent systems (Molenaar, 2022a). A hybrid system begins by off-loading some regulatory work to the AI—diagnosing knowledge gaps, recommending resources, or visualising progress—then gradually on-loads control back to the learner or the group as their self-regulatory competence grows (Nazaretsky et al., 2024; Zhang et al., 2024). Advances in GenAI amplify this potential: large language models (LLMs) can provide context-sensitive explanations, create adaptive prompts, or summarise a group’s discussion on demand (Whalen et al., 2023). At the same time, AI can automate repetitive assessment tasks and mine vast data sets for patterns invisible to teachers, freeing educators to devote their time to higher-order coaching and socio-emotional support (Cukurova, 2024). The essence of HHAIR is this dynamic handover of control, calibrated to learners’ needs and aimed at cultivating durable self- and co-regulation skills, which aims to the final goal–AI Regulation–as explained in Table 1. HHAIR emphasises the importance of human agency and oversight even when using AI assistance (Lai, 2022).
Table 1: Hybrid Human-AI Regulation Levels (Molenaar, 2022a)
Degrees of hybrid regulation Description Instrumentation Tool Examples
AI regulation AI monitors and adjusts extensively, with humans aware of its regulation Tools that are fully controlled by AI, which adapts learning processes and content while keeping learners informed of its regulatory actions to support their learning.
Co-regulation AI monitors and adjusts incrementally, and humans understand its monitoring and control Tools that collect multimodal learner data, such as facial expressions, EDA, eye-tracking, and learning traces, and use this data to adapt learning content and provide scaffolding in real-time
Shared-regulation AI monitors and suggests control actions, while humans understand monitoring and perform control Tools that access learners’ written work, use AI for analysis, and deliver feedback (Tang et al., 2024)
Self-regulation Humans monitor and initiate control independently Tools for creating plan and checking time (Winne et al., 2019)
Attending to SSRL and HHAIR is therefore crucial in contemporary education. Research shows that learners often under-utilise effective regulation strategies, while many adaptive tools silently take those decisions away, depriving students of practice (Nazaretsky et al., 2024). Thoughtfully designed AI tools and systems need to address both problems: it keeps collaboration efficient in the moment, provides personalised and timely feedback, and explicitly models the metacognitive moves for learners to internalise (Cukurova, 2024; Yan et al., 2024). However, these tools remains limited. One key reason for this gap is the complexity of designing adaptive AI systems that can dynamically share control with human learners and educators (Hao and Cukurova, 2023). Current AI-driven educational technologies are heavily focused on providing automated and fixed recommendations, often lacking mechanisms for reciprocal interaction and real-time feedback integration from learners (Roll and Winne, 2015). Most existing adaptive learning systems prioritise knowledge acquisition and task adaptation, rather than supporting SRL and metacognitive processes, which require the AI to recognise, interpret, and respond to a wide range of learner behaviours, emotions, and cognitive states dynamically (Azevedo and Gašević, 2019). Developing systems that can contextually adjust the level of intervention based on individual’s skills and autonomy remains a major technical challenge. Furthermore, most learning management systems (LMS) and AI-driven tutoring platforms lack transparent, explainable AI mechanisms to ensure that learners and teachers can understand how AI-driven decisions are made (Shin, 2021; Khosravi et al., 2022). Without clear communication of AI reasoning, users may struggle to trust or effectively engage with the system (Kizilcec, 2016). Limited interdisciplinary collaboration among AI developers, cognitive scientists, and education practitioners further contributes to the slow development of robust hybrid regulation tools (Molenaar, 2021).
The FLoRA Engine has been developed and will continue to evolve to address these challenges by harnessing expertise from multiple areas, including researchers, educators, and AI developers. Its primary objective is to bridge the gap between AI and SRL, fostering a more adaptive and personalised learning experience. To achieve this, FLoRA provides a diverse set of tools designed to support both teachers and students. Teachers can utilise these tools to design and customise course content, while also maintaining control over the degree of AI intervention in the learning process. Meanwhile, students receive personalised support to enhance their SRL skills. Additionally, learner and teacher feedbacks are continuously integrated into the system, ensuring ongoing improvements based on real-world educational interactions. Ultimately, FLoRA aspires to develop fully autonomous AI Regulation (explained in Table 1), wherein the system dynamically adjusts its support based on each learner’s progress, thereby maximizing the effectiveness of AI in education.
3 Innovative Features
This section introduces the major advancements incorporated into the FLoRA Engine, emphasizing the core innovations that distinguish FLoRA from existing tools and approaches. Particular attention is given to the most notable instrumentation tools developed, with detailed accounts of their functionalities and underlying process logic. These enhancements demonstrate the capacity of FLoRA to serve as a state-of-the-art research tool for investigating and supporting SRL within rapidly evolving educational environments shaped by the widespread adoption of GenAI technologies.
Figure 1 presents the integration of the FLoRA Engine within the Moodle platform, highlighting its user interface. Three principal zones are identified: the Navigation zone, Reading zone, and Instrumentation tools zone, with the latter being offered by the FLoRA Engine. While the Navigation and Reading zones are standard Moodle components, the Instrumentation tools zone enables comprehensive tracking and processing of all learner interactions within the web pages.
Figure 1: Overview of the FLoRA User Interface
3.1 From Learning Analytics to Personalised Scaffolds with GenAI
FLoRA Engine incorporates an analytics-based GenAI scaffolding function that integrates learning analytics and GenAI to offer personalised support for students’ SRL processes. The tool is shown in Figure 2. Its name has been updated to "Instruction Panel" to enhance user clarity. At predefined intervals, users receive scaffolding messages through this panel, providing personalized guidance during their learning process.
The mechanism that triggers this support is illustrated in Figure 3. As learners work on the given tasks, learning analytics techniques collect real-time data, capturing both internal learning conditions and SRL processes. According to Winne (2022), internal conditions comprise, for instance, a learner’s level of domain knowledge (a static condition) and their awareness of time constraints (a dynamic condition). In a time-limited learning task, a learner might initially adopt a linear approach, but as they become aware of the remaining time, they may adjust their strategies to accommodate this constraint. Consequently, time-awareness can be considered a dynamic learning condition that updates over the course of the task. In FLoRA , such changes in dynamic conditions are detected in real-time through trace data (e.g., a click on the timer to check the remaining time). Meanwhile, static conditions, such as learners’ domain knowledge levels, are determined through pre-task surveys and assessments. FLoRA automatically scores these instruments, and the resulting scores feed into the GenAI-based scaffolding tool to tailor the support provided to each learner. Each learning condition is converted into a predefined statement and added to the GenAI prompt. More details can be found in Appendix C. For example, if a student’s trace data shows they clicked on the timer, it indicates awareness of time constraints, generating “the student is aware of the time constraint.” Conversely, no click indicates unawareness, resulting in “the student is not aware of the time constraint.”
Figure 2: GenAI-based scaffolding tool
Meanwhile, SRL processes are captured in real-time representing how learners self-regulate their learning. For example, when a student working on drafting their essay, and checking task instruction and rubrics during the drafting, this sequential process will be automatically labelled as evaluation. The current FLoRA is compatible with different theoretical models informing the design of trace parser, including Winne and Hadwin’s COPES model (2022) and Zimmerman’s cyclical model of SRL (2013), which is empirically validated using think-aloud data as the reference points (Authors, 2021). Our previous lab study identified critical SRL processes that were significantly associated with learning performance at various stages of the task (Authors, 2024b). For instance, in a timed reflective writing task, students who engaged in orientation—using task instructions and marking rubrics to guide their reading within the first two minutes—scored significantly higher than those who did not (van der Graaf et al., 2023). Drawing on these findings, FLoRA can automatically trigger GenAI-powered scaffolding whenever these critical SRL processes remain undetected for a specified time period.
Figure 3: GenAI-based scaffolding trigger mechanism
As shown in Figure 3, learning conditions and critical SRL processes—collected through real-time analytics—are sent to the FLoRA system, which automatically generates a GenAI prompt using the relevant template. These prompts are then passed to GenAI to produce responses that become SRL scaffolding. The resulting feedback is delivered to students via the GenAI-based scaffolding tool, nudging them toward effective SRL processes with content tailored to their evolving needs (Figure 2). This integrated design—combining learning analytics with GenAI—offers clear advantages over rule-based scaffolding, which relies on a predetermined set of messages (Srivastava et al., 2022; van der Graaf et al., 2023). Since rule-based systems require all potential messages to be predefined, addressing multiple learning conditions simultaneously becomes unwieldy and impractical due to the sheer complexity of enumerating possible rule combinations (Backhaus et al., 2017).
Figure 4: Writing analytics tool - basic writing
3.2 From Writing Analytics to Personalised Feedback
The writing analytics tool in FLoRA is specifically designed to foster SRL within the context of information problem-solving tasks. By delivering targeted, formative feedback for revision, the tool cultivates essential SRL processes such as self-monitoring, reflection, and strategic adaptation. Its four core functions address different dimensions of writing proficiency, enabling iterative improvement and the development of critical metacognitive skills.
Figure 5: Writing analytics tool - academic writing
The first core function, the Basic Writing module (see Figure 4), focuses on identifying and correcting spelling and grammatical errors. Feedback is generated through GenAI assessments, ensuring accuracy and specificity in suggestions. By addressing foundational language issues, this module not only enhances grammatical correctness but encourages students to regularly evaluate and refine their writing—a key aspect of the SRL process of monitoring.
Figure 6: Writing analytics tool - originality
The second function, the Academic Writing module (Figure 5), supports learners in aligning their language with academic conventions. This module draws on a back-end Map data structure and NLP phrase-matching to efficiently identify issues with tone, sentence complexity, and vocabulary usage. By delivering context-specific guidance grounded in academic expectations, the tool promotes students’ ability to plan and adapt their writing style to disciplinary norms, thus supporting SRL strategies related to goal-setting, strategic planning, and self-reflection
Figure 7: Writing analytics tool - cognition classification
The Originality function (Figure 6) evaluates the uniqueness of student text by flagging any instance where more than seven consecutive words match the background reading material, leveraging an n-gram algorithm for detection. Beyond supporting academic integrity, the explicit highlighting of copied segments prompts students to monitor their use of sources, reconsider how they synthesize information, and actively paraphrase in their own words. This encourages the SRL processes of self-evaluation and strategic revision, empowering learners to internalize standards for originality.
Finally, the Cognition Classification module (Figure 7) analyses each sentence using Bloom’s taxonomy, classifying it according to cognitive level—ranging from remembering to creating—via the OpenAI davinci-002 model (Forehand, 2010; Iqbal et al., 2023, 2024). By providing real-time feedback on the depth and complexity of their arguments, students are guided to elevate their essays through higher-order thinking. This function explicitly supports SRL by encouraging ongoing evaluation of cognitive engagement and targeted revision to achieve deeper analytical and synthetic writing.
Each module delivers immediate, transparent feedback by highlighting errors or cognitive levels within the student’s text. This direct correspondence between analytic results and source text empowers students to quickly identify strengths and weaknesses and strategically address them. For example, visible error markup in Basic and Academic Writing facilitates focused self-correction, while clear identification of copied passages in the Originality function scaffolds effective revision behaviour. The Integration and Elaboration tool’s feedback encourages learners to set higher goals for cognitive engagement, aligning their efforts with the objective of continuous improvement—a fundamental principle of SRL.
By integrating these targeted feedback mechanisms, the writing analytics tool not only enhances writing quality but also actively scaffolds students’ self-regulatory cycles of goal-setting, monitoring, and adaptive revision. Consequently, FLoRA ’s analytics framework prepares learners to become independent, reflective writers capable of managing and advancing their own learning processes.
3.3 Chatbot Support with Multiple GenAI-powered Agents
With the rapid advancement of GenAI, its integration into educational contexts has become increasingly prevalent. GenAI-powered tools, such as chatbots, provide students with immediate access to information and personalized guidance, thereby offering significant potential to foster SRL. By facilitating processes such as setting learning objectives, selecting effective strategies, and reflecting on outcomes, GenAI tools can scaffold key stages of SRL. Within the FLoRA Engine, we have developed and integrated GenAI-based chatbot tools that leverage comprehensive learning contexts—including students’ prior knowledge, learning goals, and trace data logs—to systematically monitor progress and deliver real-time, actionable feedback. This continuous, contextualized support promotes metacognitive development, encouraging learners to reflect on their approaches and make adaptive adjustments throughout the learning process.
Figure 8: Multi-Agents chatbot tool – single chat window
To accommodate diverse educational scenarios, the chatbot tool can be configured and used in flexible ways. The Figure 8 employs a single chat window that supports multiple GenAI agents, each configured with distinct pre-prompts to specialise in specific domains or tasks. All agents share the conversation history, enabling learners to interact with a panel of experts within a cohesive dialogue. This collaborative environment mirrors authentic problem-solving situations, where engaging with multiple perspectives enhances the depth and breadth of feedback. By receiving collective input from specialized agents, students are better able to monitor their understanding, seek clarification, and receive nuanced guidance—thereby strengthening SRL components such as self-monitoring, strategy selection, and adaptive regulation.
Figure 9: Multi-Agents chatbot tool – separated chat window
The Figure 9 adopts a multi-window configuration, assigning each GenAI agent to a separate, independent chat interface. In this setup, conversation histories are not shared, allowing each interaction to remain focused and distinct. This design is particularly effective in simulations or role-based scenarios where compartmentalized, parallel discussions are required. For example, in a negotiation exercise, one agent may serve as the client and another as the supplier, enabling learners to engage fully with each perspective. This separation supports SRL by encouraging learners to set targeted objectives, monitor progress within specific contexts, and evaluate outcomes independently for each scenario.
Both versions of the chatbot tool offer flexible configuration options, such as customizable pre-prompts and agent avatars, allowing educators to tailor the tool to specific course requirements and learning preferences. By seamlessly integrating these advanced GenAI chatbots into the FLoRA Engine, we provide learners with dynamic, interactive, and individually responsive support mechanisms. Ultimately, these tools enhance engagement, foster critical thinking, and promote the iterative self-assessment and strategy adjustment central to effective self-regulated and hybrid human-AI regulated learning.
3.4 Collaborative Writing with Adaptive Support
To broaden the applicability of the FLoRA Engine, we have developed a collaborative writing tool designed to enhance both educational and research capabilities. This tool enables multiple users to simultaneously edit shared documents, similar to platforms like Google Docs (see Figure 10). Distinctively, FLoRA ’s collaborative tool records granular activity data—including keystrokes, revisions, and user interactions—in real time. This comprehensive data logging provides valuable insights into learners’ collaborative strategies, participation patterns, and group dynamics, thereby facilitating robust analysis of collaborative learning processes in authentic environments.
A key innovation of the FLoRA collaborative writing tool is the integration of a GenAI-powered chat interface. This embedded intelligent agent augments group communication by delivering context-sensitive feedback, answering queries, and generating relevant suggestions. The adaptive, personalized scaffolding provided by GenAI supports learners as they engage in real-time goal setting, monitor both their own and their peers’ contributions, and reflect on collaborative strategies. Consequently, students are empowered to exercise critical SRL skills such as self-monitoring, strategic revision, and adaptive planning, all within a supportive, interactive environment.
The fusion of AI and collaborative writing in FLoRA advances the concept of hybrid human–AI regulated learning. By mediating peer dialogue and offering individualized feedback, the GenAI agent fosters metacognitive development and the co-construction of knowledge. Learners actively regulate their learning by setting objectives, adjusting approaches, and evaluating outcomes, benefitting from immediate support tailored to both personal and group needs. This synergy enhances the depth and quality of collaboration, increases student engagement, and directly cultivates SRL competencies.
Figure 10: Collaborative writing tool
In addition to these innovative features, FLoRA encompasses a suite of other tools designed to support SRL across diverse educational contexts. Further details on these capabilities are provided in Appendix A.
4 FLoRA Engine Implementation in Research Studies
The FLoRA Engine has been adopted by more than 20 universities and institutions globally, supporting a substantial and diverse student population. This widespread adoption reflects its effectiveness in detecting, measuring, and facilitating SRL through advanced, research-informed tools for both educators and learners. As a result, the FLoRA Engine has become an integral resource for enhancing educational outcomes in varied instructional settings.
To illustrate its practical utility, efficacy, and adaptability, this section presents representative case studies from diverse educational contexts. These examples demonstrate how the FLoRA Engine supports a wide array of learning objectives and tasks, showcasing its versatility and positive impact on learning processes. Through an examination of its application across scenarios, we offer evidence for the Engine’s capacity to advance SRL and HHAIRL in authentic environments.
4.1 Case 1 – Comparing Human, Rule-Based, and GenAI Feedback
The first study investigated the effectiveness of different support mechanisms in enhancing English writing skills among university students who are English as a Second Language (ESL) learners. The primary objective was to assess how various forms of assistance—including GenAI, human expert guidance, and writing analytics tools—impact students’ performance in a two-stage English reading and writing task.
Participants: A total of 117 university students participated in the study conducted from July to September 2023. The participants had an average age of 22.61 years (SD = 3.39), with 70% identifying as female and 55% being undergraduates. They are from diverse disciplines and shared a common first language other than English. All were ESL learners.
Figure 11: Experimental procedure of case study 1
Research Procedure: The study was conducted in a controlled laboratory setting and followed a six-step procedure, as illustrated in Figure 11: pre-task questionnaire, Stage 1 training about the use of FLoRA , Stage 1 reading and essay writing task, Stage 2 training about the use of FLoRA , Stage 2 essay revising, and post-task questionnaire. Participants used computers to complete questionnaires, watch training videos, and engage in the learning tasks. In Stage 1, after watching a training video on how to use the learning tools provided in the learning environment, participants undertook a two-hour reading and writing task. They were required to read materials on three topics — artificial intelligence, differentiated teaching, and scaffolding teaching — and write an essay envisioning the future of education in 2035 by integrating these topics. A rubric outlining the grading criteria was provided for reference during the writing process. In Stage 2, participants watched a second training video introducing the specific support mechanisms available to their assigned experimental group. They then engaged in a one-hour revision task, aiming to improve their essays using the provided support. After completing the tasks, participants were asked to complete a post-test within one day.
Participants were randomly assigned to one of four experimental groups, each receiving distinct forms of support during the revision stage. The Control Group (CN, n = 30) received no additional support, completing the revision independently and experiencing no changes to the learning environment across both stages. The AI Group (n = 35) received assistance from a GPT-4-based chatbot tool (see Section 3.3), embedded within the user interface. This AI support was limited to task-related content through targeted pre-prompts, ensuring conversations remained relevant to the learning objectives. The Human Expert Group (HE, n = 25) received guidance from a qualified academic writing specialist. Participants in this group could interact with the expert via an instant chat tool (see Appendix A) to refine and enhance their essays. The Writing Analytics Tools Group (CL, n = 27) utilised a writing analytics toolkit (see Section 3.2), which provided automated feedback on spelling and grammar, academic style, originality, and cognition classification.
Key Research Findings: This study yielded several key findings regarding learners’ interactions with GenAI-based chatbots, human experts, and writing analytics tools. First, learners who engaged with GenAI-based chatbots exhibited more direct and pragmatic help-seeking behaviors, often bypassing metacognitive stages such as diagnosing their difficulties or evaluating the support received. This resulted in a non-linear help-seeking process focused on obtaining specific information quickly through direct, knowledge-oriented questions. In contrast, learners interacting with human experts demonstrated a more linear approach, employing indirect and evaluative questioning to establish shared understanding and maintain social dynamics (Chen et al., 2025a; Cheng et al., 2025). This divergence in questioning styles mediated the relationship between support type and academic performance: learners who asked direct questions—typically supported by GenAI—showed greater improvements in essay revisions due to efficient knowledge acquisition (Cheng et al., 2025). However, reliance on GenAI sometimes reduced metacognitive engagement, as learners tended to delegate metacognitive monitoring to the chatbot, potentially limiting deep learning and critical reflection (Chen et al., 2025a). Second, GenAI-based chatbots and human experts embodied distinct values. The chatbot fostered social comfort and autonomy by minimizing social judgment, but was limited in responsiveness and raised privacy concerns. Conversely, human experts provided personalised and responsive support but occasionally constrained learner autonomy. Learners’ value priorities varied based on personal preferences and contexts, resulting in inherent trade-offs in their learning experiences (Shen et al., 2025). Third, while the GenAI chatbot often facilitated immediate performance gains—such as higher essay scores—these benefits did not translate into sustained intrinsic motivation or long-term knowledge gains compared to human tutoring or independent work. The convenience of GenAI support at times led to “metacognitive laziness,” with learners engaging less in planning and monitoring, both vital for self-regulated learning (SRL) (Authors, 2024). Fourth, the design of formative writing analytics tools significantly influenced learners’ self-assessment processes. Tools offering high-affordance feedback, such as clear grammatical corrections, supported accurate self-assessment and substantive improvement in writing. Conversely, low-affordance tools, including those integrating Bloom’s taxonomy classifications, often hindered effective self-reflection and revision due to lack of actionable guidance or potential for confusion (Tang et al., 2024; Forehand, 2010). Finally, although learners generally preferred human feedback, preferences were shaped by the type of GenAI tool used. High-performance, interactive chatbots like ChatGPT (see Section 3.3) fostered stronger preferences for AI-supported learning, whereas less interactive or rule-based tools (see Section 3.2 and Appendix A) led learners to favour human tutors for their greater personalization and interactivity (Le et al., 2025; Cheng et al., 2025).
4.2 Case 2 – Clinical Reasoning Training
This study explores the effectiveness of using GenAI-based chatbot tool, which is adjusted with preprompt to simulate as a Virtual Standard Patient (VSP), to support medical students’ SRL and HHAIRL, with the ultimate goal of enhancing clinical reasoning skills. By simulating realistic patient interactions, the research aims to assess whether engaging with GenAI VSP improves students’ history-taking abilities, diagnostic reasoning, and overall clinical competence. Additionally, it examines the impact of a GenAI-based virtual teaching assistant on providing personalised scaffolding to support learning.
Participants: A total of 210 second-year clinical medical students from a Medical university participated in this study. All participants had completed foundational training in symptomatology and history-taking. However, many self-reported their clinical reasoning abilities as average, highlighting the need for additional skill development in this area.
Figure 12: Experimental procedure of case study 2
Research Procedure: The study consisted of five stages, illustrated in Figure 12: a pre-task survey, a training session, the main task (comprising five consecutive clinical reasoning tasks), and a post-task survey. In the pre-task survey, participants provided demographic data, previous course grades, and self-assessments of their history-taking and clinical reasoning abilities. The training session required participants to review task instructions and view a demonstration video on the FLoRA learning environment. The core of the study involved five weekly clinical reasoning tasks, each focused on a distinct medical topic including Spontaneous Pneumothorax, Stable Angina, Aortic Dissection, Acute Pulmonary Embolism, and Acute Pericarditis. Each task lasted 120 minutes. During these sessions, students interacted with the VSP via a chat interface to gather information and formulate a diagnostic conclusion. They were able to review instructions, interact with the VSP, take notes, draft, and submit diagnostic conclusions. Upon submission, students received automated GenAI-based feedback from the FLoRA system according to predefined rubrics (see Section 3.3). After completing the main tasks, participants completed a post-task survey to reassess their skills and reflect on their learning experience.
Participants were randomly assigned to either the GenAI Support group (AI Group, n = 106) or the control group (CN Group, n = 104). During modules 2–4 of the main task, the AI group interacted with both the GenAI VSP and a GenAI teaching assistant. This teaching assistant provided individualised feedback, guidance, and adaptive scaffolding to support learning. In contrast, the control group engaged only with the GenAI VSP, allowing for a direct comparison of the effects of personalised GenAI support. Both groups accessed the FLoRA platform’s suite of tools, which included the GenAI VSP chat interface powered by ChatGPT models, an auto-saving note-taking tool with data collection features, a text submission tool for diagnostic results that offered immediate feedback and performance scores (see Appendix A), and a timer to regulate task duration.
Key Research Findings: The study assessed students’ clinical reasoning skills by analysing their history-taking behaviours during interactions with VSPs. Some behaviours showed strong correlations with reasoning metrics like diagnostic accuracy, history-taking scores, and clinical knowledge test results. This is because these behaviours reflected essential components of effective clinical reasoning: organizing questions in a logical sequence aids in systematic data collection; asking specific, pathophysiologically relevant questions demonstrates a deeper understanding of disease processes; and summarising and reorganising information indicates the ability to synthesise patient data. These behaviours enhance students’ ability to accurately diagnose and understand clinical scenarios, thereby improving their performance on reasoning metrics. By integrating GenAI-driven VSPs into the FLoRA platform, the study offered scalable and interactive opportunities for students to practice history-taking in a realistic, simulated environment. This approach enabled dynamic patient consultations and provided immediate, personalised feedback, thereby enhancing students’ clinical reasoning skills (Chen et al., 2025b).
Beyond generating valuable insights in medical diagnostics, this study also demonstrates the versatility of the FLoRA platform for a wide range of experiential learning tasks. Its capacity to simulate interactive scenarios suggests potential applications in leadership training and other domains requiring experiential learning. The results highlight that integrating FLoRA with GenAI technologies can facilitate skill development and deliver personalised learning experiences across diverse educational contexts.
4.3 Case 3 – advancing analytics-based adaptive SRL scaffolding with GenAI
Figure 13: Experimental procedure of case study 3
This project investigated learners’ SRL skills in both K–12 and higher education contexts, with a focus on how learning analytics and AI can be leveraged to support robust SRL processes in the era of AI. A central innovation of this work is the integration of learning analytics with GenAI to enhance the adaptivity of SRL scaffolding. Prior research has identified the limitations of rule-based analytics, which often fail to accommodate the unique strategies and conditions of individual learners (Authors, 2025b). To overcome these limitations, this project utilised real-time learning analytics to inform GenAI models, enabling the generation of personalised and contextually relevant SRL support.
Participants: Data were collected from both secondary and higher education settings. In the K–12 context, data were gathered from 663 students across 14 schools in six countries: Brazil (86 students from one school), Colombia (33 from one school), India (93 from three schools), Australia (88 from three schools), China (213 from two schools), and the United Arab Emirates (150 from four schools). In higher education, a reflective writing task was implemented as part of an academic English writing course, with data collected from 171 undergraduate students.
Research Procedure: This project comprised four sequential learning modules. The first module, the pre-task phase, included three activities: a demographic questionnaire (collecting data such as gender, age, and first language), a domain knowledge assessment on AI applications in medicine (15 multiple-choice questions), and a learning strategy knowledge test based on the ISDIMU instrument (Bannert et al., 2021). In the latter, participants responded to hypothetical scenarios by indicating the likelihood of employing specific learning strategies; for example, identifying whether they would consult solution materials before attempting a new task.
Following the pre-task module, students entered the training phase, which was designed to familiarise them with the learning platform and its built-in tools. Instructional materials were provided in a text-and-image format to support user orientation.
The third module, the core reading and writing task, required students to compose a 200–300 word essay using information from multiple texts about AI applications in medicine. During this task, students utilised the FLoRA learning tools and were randomly assigned to one of three groups (see Figure 13). The control (CN) group completed the essay with no explicit SRL support. The Process-only (Po) group received GenAI-based scaffolds tailored to their real-time SRL processes, which were identified by analysing prior task data using time-series mapping and rain-cloud plots at seven-minute intervals (van der Graaf et al., 2023). For each interval, SRL processes exhibiting significant differences in frequency between high and low performers were selected to inform the scaffolding in the Po group. Scaffolding was triggered if critical SRL processes—such as monitoring instructions or rubric use (common among high performers)—were absent within predetermined timeframes. For example, if a learner did not review the task instructions within 14 minutes, a scaffold prompted this behaviour; if the student had already done so, the scaffold was not activated.
The Process with Condition (PwC) group received similar GenAI-generated process-based scaffolding, but with content further personalised by individual learning conditions derived from the COPES model of SRL (Winne and Hadwin, 1998; Winne, 2022). These conditions included factors such as domain knowledge and awareness of available resources. Learning analytics techniques were used to detect these real-time conditions, allowing GenAI to provide adaptive, individualised scaffolding. Details of the scaffolding mechanisms and design are provided in Section 3.1.
Upon completing the main writing task, students proceeded to the post-task module, which featured two activities: a post-test using the same 15 multiple-choice questions to assess learning gains, and a feedback literacy survey evaluating attitudes, mindsets, and behaviours related to feedback (Dawson et al., 2023).
Key Research Findings:
The design of this project has been implemented in empirical studies across both K-12 and higher education settings. In the higher education context, Authors (2025b) conducted randomised controlled trial compared the three groups as described above: CN, Po and PwC groups. Using Ordered Network Analysis (ONA), the results showed that learners in the PwC group demonstrated significantly more effective and metacognitive SRL patterns, such as frequent consultation of task instructions and rubrics, compared to the other groups. Furthermore, learner compliance (i.e., the level of adherence to the suggested processes) with the scaffolds varied, with compliers showing clear adherence to scaffold suggestions and more frequent metacognitive processes, especially in writing phases, while non-compliers exhibited more linear and less strategic learning patterns. These findings highlight the importance of incorporating both learners’ real-time SRL processes and evolving conditions in scaffold design and provide empirical evidence supporting the scalability and effectiveness of GenAI-powered adaptive scaffolding for enhancing SRL in computer-based learning environments. In another study, Authors (2025d) explored the relationship between feedback literacy and SRL processes among secondary school students across different countries. Through K-medoids clustering of self-reported feedback literacy, the study identified two distinct clusters: Proactive Feedback Engagers and Moderate Feedback Engagers. ONA revealed that Proactive Feedback Engagers initially employed less effective learning strategies but adapted significantly following scaffolded feedback, demonstrating more balanced and strategic regulation. In contrast, Moderate Feedback Engagers began with more strategic initial approaches but showed less adaptability and responsiveness to scaffolding suggestions. These findings emphasize the dynamic, context-specific nature of feedback literacy and its interplay with SRL, highlighting the need for adaptive scaffolding tailored to students’ feedback literacy profiles to enhance engagement and SRL processes.
5 Discussion
The FLoRA Engine represents a significant advancement in the development of AI-enhanced educational platforms, focusing on addressing the challenges and opportunities of hybrid Human-AI regulation in SRL. This study provides evidence of FLoRA ’s ability to both facilitate SRL processes across various contexts and domains and contribute to the broader understanding of how learners interact with AI systems within complex learning environments.
5.1 Advancing Hybrid Human-AI Regulation in Education
FLoRA demonstrates the potential of GenAI and learning analytics to advance hybrid Human-AI regulation by supporting and scaffolding learners’ SRL behaviours. The platform detects critical SRL processes, assesses real-time learning conditions and, when necessary, intervenes via personalised scaffolds designed to enhance students’ cognitive, metacognitive, and motivational regulation. This aligns closely with emerging frameworks such as Hybrid Human-AI Regulated Learning (HHAIRL) (Huang, 2024; Molenaar, 2022a), which emphasise the synergetic interplay of human agency and AI augmentations in fostering learner autonomy. For instance, the scaffolding intervention mechanisms triggered by unmet SRL benchmarks (e.g., the absence of critical planning or monitoring behaviours during timed tasks) illustrate how AI can enhance learners’ self-regulatory processes while maintaining their central role in decision-making and goal-setting.
Across all three case studies described in Section 4, the results suggest that personalised scaffolding, particularly when grounded in dynamic trace data, can promote learners’ active engagement in SRL behaviours such as goal setting, process monitoring, and self-evaluation. A notable example is the integration of GenAI personalisation in writing revision tasks, where learners in the AI-backed scaffolding groups exhibited improved essay scores, a finding consistent with prior work showing how adaptive scaffolding fosters task-relevant SRL behaviours (Järvelä et al., 2023; Edwards et al., 2025). These findings contribute to the broader literature on co-regulated SRL approaches by revealing how trace analytics combined with GenAI can dynamically adjust scaffolds to learners’ changing conditions, a capability absent in traditional rule-based systems (Huang, 2024).
However, the findings also highlight the risks of over-reliance on AI systems, echoing concerns raised in prior studies (Authors, 2024a). For instance, the case study of ESL learners using GenAI-based support for writing revealed instances of "metacognitive laziness," where learners offloaded planning and monitoring to AI systems. While this improved task-specific outcomes in the short term, it may hinder the long-term development of learners’ critical SRL abilities, a finding corroborated by research on the pitfalls of AI augmentation in learning contexts (Molenaar, 2022a).
5.2 Fostering Collaboration and Personalisation Across Contexts
The adoption of FLoRA across diverse educational domains, from reflective writing tasks to clinical reasoning and K-12 classroom settings, highlights its flexibility and scalability. Each case study shed light on the nuanced needs of learners within specific contexts, further emphasising the importance of adapting tools like FLoRA to align with disciplinary requirements. For instance, the use of the GenAI-based Virtual Standard Patientin the medical diagnostic reasoning task underscores the platform’s capability to simulate real-world problem-solving environments. The tool effectively supported learners’ organisation of diagnostic queries and improved history-taking practices, reflecting the potential for hybrid regulation to complement domain-specific pedagogical designs (Chen et al., 2025b).
The findings from Case 3 further highlight the importance of delivering personalised scaffolds not solely based on learners’ SRL processes, but also aligned with individual attributes such as learner profiles and literacy levels (e.g., feedback literacy). While prior research has primarily focused on identifying and responding to learners’ real-time SRL behaviours (Winne, 2017b; Viberg et al., 2020), our study extends this perspective by demonstrating how scaffold effectiveness increases when these supports are tailored explicitly to learner characteristics. For instance, learners’ differing levels of feedback literacy considerably influenced their receptiveness to scaffolded interventions, underscoring the need for careful alignment between scaffold design and learners’ individual profiles. GenAI offers a distinct advantage over traditional rule-based systems in this context, as it can dynamically incorporate varied learner attributes and literacy levels into scaffold design, enabling deeper personalisation beyond what is feasible through predefined logic alone. This finding supports recent theoretical arguments calling for comprehensive, learner-centred approaches in adaptive scaffolding designs within hybrid Human-AI regulated learning environments (Järvelä et al., 2023; Cukurova, 2024; Nguyen, 2025).
5.3 Implications for Future Design and Research
Second, providing learners with direct access to visual analytics presents a promising avenue for future development. Learner-facing dashboards that depict real-time SRL activities can enhance metacognitive monitoring, prompting students to reflect and make targeted adjustments to their learning strategies (Authors, 2023). Such tools can help cultivate meta-level awareness and encourage the gradual shift from external scaffold reliance to self-directed regulation.
Third, integrating Bloom’s taxonomy into FLoRA ’s writing analytics demonstrated its capacity to improve the accuracy of students’ self-assessment and metacognitive monitoring (Tang et al., 2024). However, effective use of formative feedback based on cognitive domains requires preliminary learner training (Forehand, 2010). Expanding FLoRA to include educator training modules, which complement learner-focused SRL interventions with teacher-led scaffolding, could further enhance the platform’s ability to support a diverse user base.
Finally, as highlighted in Case 3, adaptive scaffolding should consider both learners’ SRL processes and their individual characteristics, including domain knowledge and feedback literacy. Future iterations of FLoRA should leverage GenAI’s capacity to dynamically integrate real-time learner conditions, potentially using multimodal data such as eye-tracking for attention and facial expressions for confusion. This level of personalization surpasses traditional rule-based systems and supports recent theoretical calls for holistic, learner-centered analytics frameworks (Cukurova, 2024; Molenaar, 2022b), ultimately advancing hybrid human–AI regulated learning environments.
5.4 Limitations of the Study and Future Directions
While the results demonstrate FLoRA ’s capacity to support hybrid Human-AI regulation and advance SRL-focused educational research, this study has certain limitations. First, the reliance on digital trace data restricts visibility into higher-order cognitive and emotional SRL processes such as self-motivation, frustration management, or contextual sense-making (Saint et al., 2020a). Incorporating multi-modal data sources, such as physiological signals or speech analysis, may enhance the comprehensiveness of SRL detection, especially motivational and affective aspects (Raković et al., 2024), in future development and research of FLoRA . Second, the experimental tasks varied significantly in domain and structure (e.g., reflective writing versus clinical reasoning), which may partially confound comparisons of scaffolding efficacy across contexts. Future studies could standardise task designs to isolate the effect of scaffolding interventions under more controlled conditions. Furthermore, longitudinal studies are needed to examine the lasting impact of hybrid Human-AI regulation on learners’ metacognitive engagement and SRL skill development over time. Finally, the ethical implications of embedding GenAI systems into personalised learning contexts remain under-explored. For example, anonymity, privacy concerns, and data use transparency should be more systematically addressed in both the research design and platform development stages to ensure compliance with ethical standards in diverse regions and populations.
6 Conclusion
The enhanced FLoRA Engine represents a significant advancement in supporting Hybrid Human-AI Regulated Learning (HHAIRL) by providing both powerful research tools and a flexible infrastructure for further development. By integrating GenAI-driven instrumentation tools, a real-time trace parser, and adaptive scaffolding, FLoRA enables a comprehensive exploration of SRL processes and the nuanced interactions between learners and AI. The system’s ability to capture, analyse, and visualize real-time data facilitates comparisons of learning patterns with and without AI involvement, offering deeper insights into the role of AI in SRL.
Furthermore, to foster collaboration and open research, FLoRA is made available under an open-source license, allowing educators and researchers to customise, extend, and adapt its functionalities to diverse educational contexts. This flexibility ensures that FLoRA not only enhances AI-supported SRL interventions but also contributes to the ongoing advancement of HHAIRL research. By bridging the gap between automated AI-driven support and human agency, FLoRA empowers learners to actively regulate their learning while enabling researchers to refine and expand our understanding of effective human-AI collaboration in education.
Appendix A Other Features
A.1 Other Tools
In the previous version of the FLoRA Engine, a planner tool was incorporated to capture students’ planning actions during the learning process. Based on the existing studies, we found that few students chose to utilise the planner in their tasks. Feedback indicated that the planner was not sufficiently useful, and when tasks were short, creating a plan without any guidance was challenging and time-consuming. Drawing insights from a prior study (Srivastava et al., 2022), we observed that students’ learning strategies can be primarily categorised into three main types: "Read first, then write," "Read and write simultaneously," and "Write intensively while reading selectively." Inspired by these analytical results, we upgraded the planner tool (as shown in Figure 14) by providing a step-by-step guide. This enhancement aims to make the planning process more intuitive and efficient for students. With the revised planner tool, students begin by selecting their main learning strategy, such as "Read first, then write." They can then allocate time to several predefined tasks and choose their preferred reading and writing strategies. By following this guided approach, students can complete their planning process within 2–3 minutes. This streamlined method not only makes the planner more accessible and user-friendly but also encourages students to engage in effective planning, even for shorter tasks.
Figure 14: Planner tool
In addition to the planner tool, FLoRA now has a tool for mathematics (displayed in Figure 15) to further expand the application scenarios of the FLoRA Engine beyond the reading and writing tasks. This feature allows learners to edit mathematical expressions directly within the platform, facilitating the integration of complex mathematical content into learning activities. By supporting the creation and manipulation of mathematical notation, the maths tool enhances the FLoRA Engine’s functionality for STEM disciplines, such as mathematics and physics. This advancement broadens the platform’s utility, enabling it to cater to a wider range of academic subjects and learning needs.
In addition to the aforementioned tools, the FLoRA Engine also offers several complementary tools designed to support diverse learning contexts. For instance, the integrated Dictionary tool leverages the Google Translation Service to facilitate multilingual translation, enhancing accessibility for users from different linguistic backgrounds. The text submission tool allow leaners to submit text to server. The Instant Chat tool allows learners to communicate directly with teachers, fostering timely interaction and support. Furthermore, the Questionnaire tool enables the administration of surveys or questionnaires during the learning process, thereby supporting data collection and formative assessment. These supplementary tools are frequently employed as complementary resources in specialised educational settings or are utilised as control group instruments in experimental studies. Due to their straightforward interfaces and minimalist, single-window designs, these tools are not depicted here in order to conserve space.
Figure 15: Math tool
A.2 Admin Portal
The enhanced FLoRA Engine now features an upgraded Admin portal that encompasses comprehensive data management, data visualization, and instrumentation tools configuration functionalities. In the previous version, the log portal was limited to basic data access capabilities, which proved insufficient given the advancements in our instrumentation tools. With the current version capturing richer and more detailed data, solely supporting data downloads no longer meets the evolving needs of researchers.
Figure 16: Admin portal - data management
To address this limitation,the data management portal’s functionalities (displayed in Figure 16) have been significantly enhanced. Researchers can now not only access and download data but also view essential metadata such as the total number of participants in a given experiment, the average number of trace data records per participant, and the number of data entries generated by different tools. This additional information provides researchers with a clearer overview of their datasets, facilitating more informed analyses. The data search process has also been improved to support fuzzy matching using account names or experiment names. This feature streamlines the data retrieval process, making it more efficient for researchers to locate specific datasets relevant to their studies.
Figure 17: Admin portal in FLoRA – visual data analytics for researchers
In terms of visual learning analytics, FLoRA has various graphical representations to depict data from multiple perspectives. For instance, Figure 17 illustrates the proportion of time different students spend on various learning actions within a single task. Such visualizations enable researchers to gain deeper insights into students’ learning behaviours and patterns, thereby supporting more nuanced interpretations of the data.
Figure 18: Admin portal in in FLoRA – configuration of tools for different study settings
Finally, to enhance the usability of the FLoRA Engine for researchers, FLoRA now has instrumentation tools configuration utility (as shown in Figure 18). This tool assists researchers in selecting and setting up the specific tools they wish to employ in their studies. By simplifying the customization process, we aim to empower researchers to tailor the platform to their unique research needs effectively.
Appendix B Architecture
The enhanced architecture of the FLoRA Engine is illustrated in Figure 19. In this upgraded version, the FLoRA Engine adopts a microservice architecture, wherein each service is responsible for specific tasks, promoting modularity and scalability. As shown in Figure 19, the entire system is segmented into four primary components: the Instrumentation Tools, the Control Service, the OmniAI Service, and the Admin Portal.
Figure 19: FLoRA Architecture
The Instrumentation Tools, developed using JavaScript, serve as the frontend interface and are designed to integrate seamlessly into learning environments such as Moodle. These tools capture user interactions and learning behaviours, collecting a different of data streams including mouse movements, clicks, and interactions with various areas or tools. This data reflects learners’ engagement and the specific learning actions they undertake during the learning process.
Once collected, the data is transmitted to the Control Service via middleware technologies including Kafka, Redis, ElasticSearch, which facilitate efficient data streaming and handling. The Control Service persists the raw data into databases and performs automatic analyses to assess learners’ progress. For instance, it can identify learners’ engagement in SRL processes. Based on the analysis results, the system provides personalised support, such as scaffolding, to enhance the learning experience and address individual learner needs.
The Control Service is built upon the robust Java frameworks Spring Boot and Spring Cloud. These frameworks are widely recognised for facilitating the development of scalable, secure, and efficient microservices. Spring Boot streamlines the creation of production-ready applications by minimizing boilerplate code and configuration efforts, allowing for rapid development and deployment. Spring Cloud extends these capabilities by providing tools for distributed systems, such as service discovery, configuration management, and circuit breakers. The combined use of Spring Boot and Spring Cloud enhances the Control Service’s ability to handle large-scale user access and high concurrency, ensuring reliability and performance. Notably, industry-leading companies like Netflix and Alibaba utilise these frameworks to support their microservices architectures, handling vast amounts of data and client requests efficiently.
During students’ learning activities, the OmniAI Service manages all machine learning-related processing. This includes sending requests to various large language model (LLM) APIs, such as OpenAI’s ChatGPT and Google’s language models, as well as conducting text analyses using machine learning models like Sentence Transformers. The OmniAI Service performs these advanced computational tasks to provide intelligent responses, feedback, and analyses that enhance the learning experience. Developed using the Python Flask framework, the OmniAI Service benefits from Flask’s lightweight yet powerful features. Flask is well-suited for building microservices and handling machine learning tasks due to its simplicity and flexibility. Its minimalistic design allows developers to implement necessary functionalities without unnecessary overhead, while its extensive ecosystem and strong community support facilitate the incorporation of various extensions and tools. By utilizing Flask, the OmniAI Service efficiently manages the computational demands of machine learning operations while maintaining scalability and performance.
The Admin Portal is responsible for overseeing data management, data visualization, and the configuration of instrumentation tools, with backend processing managed by the Control Service. In previous versions, the log portal’s functionality was limited to basic data access, primarily allowing users to download data. With significant enhancements in the current version’s instrumentation tools, the data captured has become richer and more complex. As a result, merely supporting data downloads is no longer sufficient to meet researchers’ needs. Therefore, we have significantly enhanced the functionality of the data management portal (displayed in Figure 16).
The upgraded data management portal enables researchers not only to access and download data but also to view essential metadata and statistics. Researchers can now examine basic information such as the total number of participants in a study, the average number of trace data entries per participant, and the number of data entries generated by different tools. This additional information provides valuable insights into the data’s scope and composition, facilitating more informed analyses. To streamline the data retrieval process, the portal supports fuzzy matching in data searches using account names or experiment titles, enhancing usability and saving time.
For data visualization, the portal offers various graphical representations to visualize data from multiple perspectives. For instance, Figure 17 illustrates the proportion of time different students spend on various learning actions within a task. Such visualizations enable researchers to gain deeper insights into students’ learning behaviours and patterns, supporting more nuanced interpretations of the data. This enhanced visualization capability is crucial for identifying trends and informing instructional strategies.
To facilitate easier use of FLoRA by researchers, FLoRA has an instrumentation tools configuration utility (as shown in Figure 18). This utility assists researchers in selecting and configuring the specific tools they wish to employ in their studies, simplifying the customization process. By making the configuration of instrumentation tools more accessible, we empower researchers to tailor the platform to their unique research needs effectively.
The Admin Portal is built using React, a popular JavaScript library for building user interfaces. React’s component-based architecture allows for the creation of reusable UI components, leading to efficient development and maintenance. Its virtual DOM implementation enhances performance by minimizing direct manipulation of the actual DOM, resulting in faster rendering and a smoother user experience. Moreover, React facilitates the adoption of a front-end and back-end separation architecture (also known as a decoupled or headless architecture). This approach separates the user interface (front-end) from the server-side logic and data management (back-end), offering several benefits.
The separation of front-end and back-end concerns improves scalability, as each layer can be scaled independently to accommodate increased user loads or data volumes. It enhances flexibility and maintainability, allowing developers to update the front-end without affecting back-end services and vice versa. For example, if the number of users increases significantly, additional server instances can be allocated to the Control Service to handle increased data processing, while the front-end can be optimised separately to manage higher user traffic. This modularity is essential for maintaining performance and reliability in a system designed to support large-scale educational environments.
All middleware used in the FLoRA Engine, including Redis, ElasticSearch, and Kafka, are capable of cluster deployment to handle high-throughput data processing. Redis provides fast in-memory data storage, improving the efficiency of data retrieval operations. ElasticSearch offers powerful search and analytics capabilities, enabling quick querying of large datasets. Kafka facilitates real-time data streaming and messaging, ensuring efficient communication between services. By leveraging these robust middleware solutions, the FLoRA Engine effectively manages the substantial amounts of data generated by users, maintaining performance and reliability even under heavy loads.
Appendix C GPT-based Scaffolding Triggering Logic
Condition Type Learning Condition How the Conditions were Captured
Dynamic conditions Strategic plan-making action detection: SAVE_PLANNER
Awareness of the time constraint action detection: TIMER
Awareness of the available instrumentation tools action detection: TRY_OUT_TOOLS
Awareness of the available reading material action detection: PAGE_NAVIGATION
Awareness of the task requirement action detection: TASK_REQUIREMENT
Awareness of the available marking rubric action detection: RUBRIC
Static conditions Level of knowledge of learning strategies ISDIMU (Bannert et al., 2021) questionnaire score
Level of prior knowledge Pre-test score
Table 2: Learners’ Learning Conditions (Authors, 2025b)
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to check grammar of the paper. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.
[…]